# hyperExp head to column names, separator = comma, centered tabular = —c—c—c—c—c—c—c—c—c—c—, table head =     FW   FW EleSym   AGM   AGM EleSym Error Rel. time S(%) Rel. time S(%) Rel. time S(%) Rel. time S(%) , late after line =

### v3

@@ 3.3 Limitations and further discussion
The first is that it is not known how large the class of isometric hyperbolic polynomials is and even hyperbolic polynomials associated to polyhedral cones may fail to be isometric. In a sense, this is not surprising because the same cone can be generated by different hyperbolic polynomials, e.g., +$\mathbb{R}^{3}_{+}$ also satisfies +$\mathbb{R}^{3}_{+}=\Lambda(p,(1,1,1))$ for $p(x_{1},x_{2},x_{3})=x_{1}^{2}x_{2}^{2}x_{3}^{2}$. An example of non-isometric polynomial is given in [BGLS01, Example 5.2], however it is less than ideal because there there are “redundancies” in the description of the underlying cone [^1] Example 5.2 in [BGLS01] corresponds to the restriction of the polynomial $p(x_{1},x_{2},x_{3})=x_{1}x_{2}x_{3}$ to a certain two-dimensional space. Every pointed two-dimensional closed convex cone is isomorphic to +$\mathbb{R}^{2}_{+}$, which only requires a degree $2$ hyperbolic polynomial. So, there is indeed a redundancy in the expression provided for Example 5.2 in [BGLS01]. .
@@ 4 A FW algorithm for strongly convex optimization over regular cones
$\min_{x\in\mathbb{R}^{n}}$ $\quad f(x)$ (P)
$\mathrm{s.t.}$ $\quad Tx+b\in{\mathcal{K}}$
$\min_{y\in\mathbb{R}^{m}}$ $\quad(f^{*}\circ T^{*})(y)+\langle b,y\rangle$ (D)
$\mathrm{s.t.}$ $\quad y\in{\mathcal{K}}^{*},$
@@
[…]
@@ 3. $(iii)$
$T^{*}y\in\mathrm{ri}\,(\mathrm{dom}\>f^{*})=\mathbb{R}^{n}$ (4.3)
Under (4.3), we can invoke an appropriate version of Fenchel’s duality theorem (e.g., [rockafellar, Corollary 31.2.1]) to conclude that (P) and (4.2) have the same optimal value and $p^{*}$ is attained if finite. In particular if (P) is feasible, then, since (4.2) is feasible as well, $p^{*}$ is attained. Recalling that the optimal value of (4.2) is $-d^{*}$, we conclude that $p^{*}+d^{*}=0$. This proves items $(i)$ and $(ii)$.
@@
Similarly, (4.5) is equivalent to $-{y}_{\mathrm{opt}}\in\partial{\delta_{{\mathcal{K}}-b}}(T{x}_{\mathrm{opt}})$. An important consequence of (4.6) is that the unique optimal solution of (4.1) can be derived from any optimal solution to the dual problem (D).
@@ 4.1 Overcoming the challenges of constructing a Frank-Wolfe based method
The first issue is to decide which side of the problem to solve: (P) or (D). We are mainly interested in (P), but solving (D) would also be enough in view of the fact that an optimal solution to (P) can be obtained via (4.6).
That said, in order for the FW method to be efficient, we need to be efficient in solving the subproblem “$\operatorname*{arg\,min}_{x\in C}\langle\nabla f(x_{k}),x\rangle$” that appears in Algorithm 1. So the decision to apply FW to either (P) or (D) depends on which side is more likely to afford an easily solvable subproblem.
After carefully examining this question, it seemed to us that solving (D) is more promising because the FW subproblem in this case can be connected to the generalized minimum eigenvalue problem over a regular cone, as we shall discuss shortly.
Having settled for (D), we discuss in the sequel two outstanding issues.
@@
First, since 00 is always contained in (4.8), the set in (4.8) is non-empty. It is also closed and convex because ${\mathcal{K}}^{*}$ is a closed convex cone. Next, we prove the boundedness of (4.8) by checking that the recession cone of (4.8) is trivial. Since the set in (4.8) contains zero, its recession cone coincides with the set of elements $y$ such that $\lambda y$ belong to (4.8) for all $\lambda\geq 0$, e.g., see [rockafellar, Theorem 8.3]. All such $y$ must then belong to ${\mathcal{K}}^{*}$ and satisfy $\langle e,y\rangle\leq 0$. However, since $e$ is an interior point of ${\mathcal{K}}$, we have $\langle e,y\rangle=0$ which forces $y=0$. Finally, (4.8) contains at least one optimal solution of (D) by Assumption 1. ∎
@@
where, we use $y_{k}$ in place of $x_{k}$ since we are working from the dual side. In what follows, it will be helpful to define
@@
$\begin{array}[]{ll}\displaystyle\min_{s\in\mathbb{R}^{m},\alpha\in\mathbb{R}}&\langle Tx_{k}+b,s\rangle\\ \mathrm{s.t.}&\langle e,s\rangle+\alpha=c_{D}\\ &s\in{{\mathcal{K}}^{*}},\alpha\in\mathbb{R}_{+}\end{array}$ (4.11)
@@ 3. $(iii)$
Item $(i)$ First we check that (4.11) satisfies Slater’s condition. ${\mathcal{K}}$ is a regular cone by assumption, so $\mathrm{int}\,{\mathcal{K}}^{*}$ is not empty. Let $s\in\mathrm{int}\,{\mathcal{K}}^{*}$. If $\langle e,s\rangle=0$, then since $e$ is an interior point of ${\mathcal{K}}$, we have $s=0$, which is impossible since $0\in\mathrm{int}\,{\mathcal{K}}^{*}$ implies ${\mathcal{K}}^{*}=\mathbb{R}^{m}$. Therefore, $s$ satisfies $\langle e,s\rangle>0$. Let
@@
Item $(ii)$ We divide the proof in two cases.
@@
Item $(iii)$ If $t_{\mathrm{opt}}=\min(0,\lambda_{\mathrm{min}}(Tx_{k}+b))=0$, then $\lambda_{\mathrm{min}}(Tx_{k}+b)\geq 0$ which implies that $Tx_{k}+b\in{\mathcal{K}}$, by Proposition 2.2. In this case, we already verified in Case (a) of item $(ii)$ that $(0,c_{D})$ is an optimal solution to (4.11). Suppose that $t_{\mathrm{opt}}<0$. Since both (4.11) and (4.12) satisfy Slater’s condition, the following conditions from classical conic linear programming duality theory are necessary and sufficient for optimality
@@
Therefore, if $t_{\mathrm{opt}}<0$, then complementary slackness (i.e., (4.16)) and $\alpha\geq 0$ implies that the optimal $\alpha^{*}$ in (4.11) is 00. In particular, the $s\in{\mathcal{K}}^{*}$ that are optimal for (4.11) are exactly the ones that satisfy $\langle e,s\rangle=c_{D}$ and $\langle s,z_{\mathrm{opt}}\rangle=0$. ∎
@@
In view of Theorem 4.3, the final piece we need to complete our discussion is a method to find a nonzero vector in $\mathcal{F}_{{z_{\mathrm{opt}}}}^{\Delta}$. Unfortunately, if ${\mathcal{K}}=\Lambda(p,e)$ is an arbitrary hyperbolicity cone, it seems nontrivial to obtain a formula for $\mathcal{F}_{z}^{\Delta}$ given an arbitrary $z\in\Lambda(p,e)$. That said, a specific nonzero conjugate vector can be obtained easily by using the following proposition. Here, we recall that $\operatorname{mult}(z)$ denotes the number of zero eigenvalues of $z$ and $p^{(i)}$ is the $i$-th directional derivative of $p$ along a fixed hyperbolic direction, see Section 2.1.
@@
$\langle\nabla p(z),y\rangle\geq 0,\forall y\in\Lambda\quad\text{ and }\quad\langle\nabla p(z),z\rangle=0.$
@@
holds. Therefore, $x\in(\partial^{1}\Lambda^{(r-1)})\cap\Lambda^{(r-2)}$. Letting $\bar{\Lambda}\coloneqq\Lambda^{(r-1)}$, the multiplicity of 00 as an eigenvalue of $z$ with respect to $p^{(r-1)}$ is one, so we can apply the previous case to $\bar{\Lambda}$ and $p^{(r-1)}$ to conclude that $\nabla p^{(r-1)}(z)\neq 0$ holds and $\nabla p^{(r-1)}(z)$ belongs to the conjugate face of $\Lambda^{(r-1)}$ at $z$. However, since $\Lambda\subseteq\Lambda^{(r-1)}$, we have ${(\Lambda^{(r-1)})}^{*}\subseteq\Lambda^{*}$. This implies that $\nabla p^{(r-1)}(z)$ belongs to the conjugate face of $\Lambda$ at $x$. ∎
@@
Before we move on, we remark there are many useful cones ${\mathcal{K}}$ for which $\mathcal{F}_{z}^{\Delta}$ is completely known, given a particular $z\in{\mathcal{K}}$. Here we list a few. For symmetric cones (this includes the case of positive semidefinite matrices and second order cones), formulae are given in [FB06, Theorem 2] and [L17, Section 4.1.1]. For $p$-cones with $p\in(1,\infty)$, see [LLP21, Section 4.1]. Power cones and exponential cones are linearly isomorphic to their dual cones under the Euclidean inner product, so with some adjustments, the formulae discussed in [LLLP22, Section 3.1] and [LLP22, Section 4.1], respectively, can also be used to determine conjugate faces for these two cones.
@@ 4.2 The proposed method and its convergence analysis
$\lim_{k\to\infty}{\mathrm{dist}\,}(y_{k},{Y}_{\mathrm{opt}})=0$ (4.18)
@@
$f(x_{k})-f({x}_{\mathrm{opt}})\leq L\sqrt{\frac{2\|T\|_{\mathrm{op}}^{2}}{\mu\lambda_{\min}(T^{*}T)}}\sqrt{h(y_{k})-d^{*}},$
@@ 4.3 Practical considerations
$\frac{1}{i!}p^{(i)}(x)=\frac{1}{d}\sum_{j=1}^{d}\omega^{-ij}p(x+\omega^{j}e)\quad(i=1,\cdots,d-1)$
@@
At the end of Section 9 in [renegar2004hyperbolic], equations for $p^{(i)}(x)$ and $\nabla p^{(i)}(x)$ are given in which $i$ appears in place of $-i$ in the “$\omega^{-ij}$” term. We believe this is a typo, which can be verified by considering $p(x_{1},x_{2},x_{3})=x_{1}x_{2}x_{3}$, $e=(1,1,1)$ and computing $p^{(1)}(x_{1},x_{2},x_{3})$ which is $x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3}$. That said, the inverse of the Vandermonde matrix given in Section 9 is correct.
@@ 4.4 Minimizing a positive definite quadratic function
$=\frac{1}{2}\langle{x}_{\mathrm{opt}},Q{x}_{\mathrm{opt}}\rangle+\langle c,{x}_{\mathrm{opt}}\rangle$
@@
$=\frac{1}{2}\langle T^{*}{y}_{\mathrm{opt}},Q^{-1}T^{*}{y}_{\mathrm{opt}}\rangle-\frac{1}{2}\langle c,Q^{-1}c\rangle$
@@ 2. $(2)$
Error      In the tables, this describes the target relative error with respect to the reference solver, which is DDS in Section 5.1 and Mosek in Section LABEL:sec:experiment_proj_p. See also explanation in Section 5.1.1 and the discussion around (5.6).
FW      Algorithm 2.
FW EleSym      Algorithm 2 using a divide-and-conquer approach to evaluate elementary symmetric polynomials and their gradients.
@@
Rel. Time      Relative time to reach the error target with respect to the reference solver. See explanation in Section 5.1.1 or Table LABEL:table:hyper_high_deg.
S(%)      Success rate, see explanation in Section 5.1.1.
@@ 5.1 Projection onto derivative relaxations
In particular, the $k$-th derivative relaxation of +$\mathbb{R}^{n}_{+}$ (see Section 2.1), denoted by $\mathbb{R}_{+}^{n,(k)}$ satisfies
@@
where $\mathbb{R}_{+}^{n,(k)}$ is the $k$-th derivative cone of +$\mathbb{R}^{n}_{+}$ along $e=\bm{1}_{n}$ and $c$ is the vector we wish to project onto $\mathbb{R}_{+}^{n,(k)}$. Since the objective function of (5.1) is a positive definite quadratic function, (5.1) is implementable as discussed in Section 4.4. The derivative relaxations +$\mathbb{R}^{n,(k)}_{+}$ for $30<k<n-$ 3, $n\geq 4$ are non-polyhedral and there are no known formulae for their orthogonal projections.
@@
(5.1)$\iff\left\{\begin{array}[]{ll}\displaystyle\min_{x,y}&y\\ \mathrm{s.t.}&y\geq\|x-c\|\\ &x\in\mathbb{R}_{+}^{n,(k)}\end{array}\right.\iff\left\{\begin{array}[]{ll}\displaystyle\min_{x,y,z}&y\\ \mathrm{s.t.}&y\geq\|z\|\\ &z=x-c\\ &x\in\mathbb{R}_{+}^{n,(k)}\end{array}\right.$ (5.3)

### v4

@@ 3.3 Limitations and further discussion
The first is that it is not known how large the class of isometric hyperbolic polynomials is and even hyperbolic polynomials associated to polyhedral cones may fail to be isometric. In a sense, this is not surprising because the same cone can be generated by different hyperbolic polynomials, e.g., $\mathbb{R}^{3}_{+}$ also satisfies $\mathbb{R}^{3}_{+}=\Lambda(p,(1,1,1))$ for $p(x_{1},x_{2},x_{3})=x_{1}^{2}x_{2}^{2}x_{3}^{2}$. An example of non-isometric polynomial is given in [BGLS01, Example 5.2], however it is less than ideal because there there are “redundancies” in the description of the underlying cone [^1] Example 5.2 in [BGLS01] corresponds to the restriction of the polynomial $p(x_{1},x_{2},x_{3})=x_{1}x_{2}x_{3}$ to a certain two-dimensional space. Every pointed two-dimensional closed convex cone is isomorphic to $\mathbb{R}^{2}_{+}$, which only requires a degree $2$ hyperbolic polynomial. So, there is indeed a redundancy in the expression provided for Example 5.2 in [BGLS01]. .
@@ 4 A FW algorithm for strongly convex optimization over regular cones
$\min_{x\in\mathbb{R}^{n}}$ $f(x)$ (P)
$\mathrm{s.t.}$ $Tx+b\in{\mathcal{K}}$
$\min_{y\in\mathbb{R}^{m}}$ $(f^{*}\circ T^{*})(y)+\langle b,y\rangle$ (D)
$\mathrm{s.t.}$ $y\in{\mathcal{K}}^{*},$
@@
Before we move on further, some remarks are in order. We suppose that $\mathbb{R}^{m}$ is equipped with some arbitrary inner product $\langle\cdot,\cdot\rangle$ for which the corresponding norm is given by $\left\|\cdot\right\|$. A similar remark applies to $\mathbb{R}^{n}$ and while $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ may have different inner products, for simplicity we will use the same symbols $\langle\cdot,\cdot\rangle$, $\left\|\cdot\right\|$ to denote the inner product and the norm in both spaces. Furthermore, the conjugate $f^{*}$, the adjoint $T^{*}$ and the dual cone $\mathcal{K}$$^{*}$ are, of course, computed with respect to the inner product in the spaces that they are defined. That said, in contrast to Section 3, one important point is that, even if $\mathcal{K}$} is a hyperbolicity cone, we will not require that $\left\|\cdot\right\|$ be the norm induced by underlying hyperbolic polynomial $p$.
As mentioned in the Section 1, (P) contains as a particular case the projection problem, since we can take $b=0$, let $T$ be the identity map and $f(x)\coloneqq\left\|x-x_{0}\right\|^{2}$, where $x_{0}$ is fixed and $\left\|\cdot\right\|$ is the norm induced by some inner product on $\mathbb{R}^{n}$. However, (P) will also allow us to handle more general quadratic minimization problems, as we shall see in Section 4.4.
@@ 3. $(iii)$
$T^{*}y\in\mathrm{ri}\,(\mathrm{dom}\>f^{*})=\mathbb{R}^{n}.$ (4.3)
Under (4.3), we can invoke an appropriate version of Fenchel’s duality theorem (e.g., [rockafellar, Corollary 31.2.1]) to conclude that (P) and (4.2) have the same optimal value and $p^{*}$ is attained if finite. In particular if (P) is feasible, then, since (4.2) is feasible as well, $p^{*}$ is attained. Recalling that the optimal value of (4.2) is $d^{*}$, we conclude that $p^{*}+d^{*}=0$. This proves items $(i)$ and $(ii)$.
@@
Similarly, (4.5) is equivalent to ${y}_{\mathrm{opt}}\in\partial{\delta_{{\mathcal{K}}-b}}(T{x}_{\mathrm{opt}})$. An important consequence of (4.6) is that the unique optimal solution of (4.1) can be derived from any optimal solution to the dual problem (D).
@@ 4.1 Overcoming the challenges of constructing a Frank-Wolfe based method
The first issue is to decide which side of the problem to solve: (P) or (D). Overall, we decided to solve (D) because the FW subproblem in this case can be connected to the generalized minimum eigenvalue problem over a regular cone, as we shall discuss shortly. Furthermore, an optimal solution to (P) can be obtained via (4.6), provided that Slater’s condition for (P) is satisfied as in Proposition 4.1. Having settled for (D), we discuss in the sequel two outstanding issues.
@@
First, since 0 is always contained in (4.8), the set in (4.8) is non-empty. It is also closed and convex because $\mathcal{K}$$^{*}$ is a closed convex cone. Next, we prove the boundedness of (4.8) by checking that the recession cone of (4.8) is trivial. Since the set in (4.8) contains zero, its recession cone coincides with the set of elements $y$ such that $\lambda y$ belong to (4.8) for all $\lambda\geq 0$, e.g., see [rockafellar, Theorem 8.3]. All such $y$ must then belong to $\mathcal{K}$$^{*}$ and satisfy $\langle e,y\rangle\leq 0$. However, since $e$ is an interior point of $\mathcal{K}$}, we have $\langle e,y\rangle=0$ which forces $y=0$. Finally, (4.8) contains at least one optimal solution of (D) by Assumption 1. ∎
@@
where we use $y_{k}$ in place of $x_{k}$ since we are working from the dual side. In what follows, it will be helpful to define
@@
$\begin{array}[]{ll}\displaystyle\min_{s\in\mathbb{R}^{m},\alpha\in\mathbb{R}}&\langle Tx_{k}+b,s\rangle\\ \mathrm{s.t.}&\langle e,s\rangle+\alpha=c_{D}\\ &s\in{{\mathcal{K}}^{*}},\alpha\in\mathbb{R}_{+}.\end{array}$ (4.11)
@@ 3. $(iii)$
Item (i)(i)First we check that (4.11) satisfies Slater’s condition. $\mathcal{K}$} is a regular cone by assumption, so $\mathrm{int}\,{\mathcal{K}}^{*}$ is not empty. Let $s\in\mathrm{int}\,{\mathcal{K}}^{*}$. If $\langle e,s\rangle=0$, then since $e$ is an interior point of $\mathcal{K}$}, we have $s=0$, which is impossible since $0\in\mathrm{int}\,{\mathcal{K}}^{*}$ implies ${\mathcal{K}}^{*}=\mathbb{R}^{m}$. Therefore, $s$ satisfies $\langle e,s\rangle>0$. Let
@@
Item (ii)(ii)We divide the proof in two cases.
@@
Item (iii)(iii)If $t_{\mathrm{opt}}=\min(0,\lambda_{\mathrm{min}}(Tx_{k}+b))=0$, then $\lambda_{\mathrm{min}}(Tx_{k}+b)\geq 0$ which implies that $Tx_{k}+b\in{\mathcal{K}}$, by Proposition 2.2. In this case, we already verified in Case (a) of item $(ii)$ that $(0,c_{D})$ is an optimal solution to (4.11). Suppose that $t_{\mathrm{opt}}<0$. Since both (4.11) and (4.12) satisfy Slater’s condition, the following conditions from classical conic linear programming duality theory are necessary and sufficient for optimality
@@
Therefore, if $t_{\mathrm{opt}}<0$, then complementary slackness (i.e., (4.16)) and $\alpha\geq 0$ implies that the optimal $\alpha^{*}$ in (4.11) is 0. In particular, the $s\in{\mathcal{K}}^{*}$ that are optimal for (4.11) are exactly the ones that satisfy $\langle e,s\rangle=c_{D}$ and $\langle s,z_{\mathrm{opt}}\rangle=0$. ∎
@@
In view of Theorem 4.3, the final piece we need to complete our discussion is a method to find a nonzero vector in $\mathcal{F}_{{z_{\mathrm{opt}}}}^{\Delta}$. Unfortunately, this is a task that is very specific to the cone at hand, so now we take a look at some particular cases.
First, if ${\mathcal{K}}=\Lambda(p,e)$ is an arbitrary hyperbolicity cone, it seems nontrivial to obtain a formula for $\mathcal{F}_{z}^{\Delta}$ given an arbitrary $z\in\Lambda(p,e)$. That said, a specific nonzero conjugate vector can be obtained easily by using the following proposition. Here, we recall that $\operatorname{mult}(z)$ denotes the number of zero eigenvalues of $z$ and $p^{(i)}$ is the $i$-th directional derivative of $p$ along a fixed hyperbolic direction, see Section 2.1.
@@
$\langle\nabla p(z),y\rangle\geq 0,\forall y\in\Lambda\qquad\text{ and }\qquad\langle\nabla p(z),z\rangle=0.$
@@
holds. Therefore, $x\in(\partial^{1}\Lambda^{(r-1)})\cap\Lambda^{(r-2)}$. Letting $\bar{\Lambda}\coloneqq\Lambda^{(r-1)}$, the multiplicity of 0 as an eigenvalue of $z$ with respect to $p^{(r-1)}$ is one, so we can apply the previous case to $\bar{\Lambda}$ and $p^{(r-1)}$ to conclude that $\nabla p^{(r-1)}(z)\neq 0$ holds and $\nabla p^{(r-1)}(z)$ belongs to the conjugate face of $\Lambda^{(r-1)}$ at $z$. However, since $\Lambda\subseteq\Lambda^{(r-1)}$, we have ${(\Lambda^{(r-1)})}^{*}\subseteq\Lambda^{*}$. This implies that $\nabla p^{(r-1)}(z)$ belongs to the conjugate face of $\Lambda$ at $x$. ∎
@@
Before we move on, we mention other useful cones $\mathcal{K}$} for which $\mathcal{F}_{z}^{\Delta}$ is completely known, given a particular $z\in{\mathcal{K}}$. For symmetric cones (this includes the case of positive semidefinite matrices and second order cones), formulae are given in [FB06, Theorem 2] and [L17, Section 4.1.1]. For $p$-cones with $p\in(1,\infty)$, see [LLP21, Section 4.1]. Power cones and exponential cones are linearly isomorphic to their dual cones under the Euclidean inner product, so with some adjustments, the formulae discussed in [LLLP22, Section 3.1] and [LLP22, Section 4.1], respectively, can also be used to determine conjugate faces for these two cones.
@@ 4.2 The proposed method and its convergence analysis
$\lim_{k\to\infty}{\mathrm{dist}\,}(y_{k},{Y}_{\mathrm{opt}})=0.$ (4.18)
@@
$f(x_{k})-f({x}_{\mathrm{opt}})\leq L\sqrt{\frac{2\|T\|_{\mathrm{op}}^{2}}{\mu\lambda_{\min}(T^{*}T)}}\sqrt{h(y_{k})-d^{*}}.$
@@ 4.3 Practical considerations
$\frac{1}{i!}p^{(i)}(x)=\frac{1}{d}\sum_{j=1}^{d}\omega^{-ij}p(x+\omega^{j}e)\qquad(i=1,\cdots,d-1)$
@@
At the end of Section 9 in [renegar2004hyperbolic], equations for $p^{(i)}(x)$ and $\nabla p^{(i)}(x)$ are given in which $i$ appears in place of i-i in the $\omega^{-ij}$” term. We believe this is a typo, which can be verified by considering $p(x_{1},x_{2},x_{3})=x_{1}x_{2}x_{3}$, $e=(1,1,1)$ and computing $p^{(1)}(x_{1},x_{2},x_{3})$ which is $x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3}$. That said, the inverse of the Vandermonde matrix given in Section 9 is correct.
@@ 4.4 Minimizing a positive definite quadratic function
$=\frac{1}{2}\langle{x}_{\mathrm{opt}},Q{x}_{\mathrm{opt}}\rangle+\langle c,{x}_{\mathrm{opt}}\rangle.$
@@
$=\frac{1}{2}\langle T^{*}{y}_{\mathrm{opt}},Q^{-1}T^{*}{y}_{\mathrm{opt}}\rangle-\frac{1}{2}\langle c,Q^{-1}c\rangle.$
@@ 2. $(2)$
Error      In the tables, this describes the target relative error with respect to the reference solver, which is DDS in Section  5.1 and Mosek in Section  LABEL:sec:experiment_proj_p. See also explanation in Section  5.1.1 and the discussion around (5.6).
FW      Algorithm  2.
FW EleSym      Algorithm  2 using a divide-and-conquer approach to evaluate elementary symmetric polynomials and their gradients.
@@
Rel. Time      Average relative time to reach the error target with respect to the reference solver. Values close to 0 are better. See explanation in Section  5.1.1.
S(%)      Success rate, see explanation in Section  5.1.1.
@@ 5.1 Projection onto derivative relaxations
In particular, the $k$-th derivative relaxation of $\mathbb{R}^{n}_{+}$ (see Section 2.1), denoted by $\mathbb{R}_{+}^{n,(k)}$ satisfies
@@
where $\mathbb{R}_{+}^{n,(k)}$ is the $k$-th derivative cone of $\mathbb{R}^{n}_{+}$ along $e=\bm{1}_{n}$ and $c$ is the vector we wish to project onto $\mathbb{R}_{+}^{n,(k)}$. Since the objective function of (5.1) is a positive definite quadratic function, (5.1) is implementable as discussed in Section 4.4. The derivative relaxations $\mathbb{R}^{n,(k)}_{+}$ for 0<k<n30<k<n-3, $n\geq 4$ are non-polyhedral and there are no known formulae for their orthogonal projections.
@@
[…]